32, Give 1000 mL D5½NS q6h. Drop factor is 15 gtt/mL.
Answer #1

Number of drops per minute = total number of fluids/total number of hours *60. * drop factor

= 1000/6*60. * 15

= 1000 / 360 * 15

= 2.77 * 15

= 41.55

= 42 ( approximately )

We have to set 42 drops per minute to infuse 1000 ml over 6 hours .